### v2

# Borel 1 type mappings and the respective equi-families

Abstract.
We investigate classes of functions from a topological space to a metric space that are related to those of Borel class 1. Following the idea defining an equi-Baire 1 family (due to Lecomte) we define the respective equi-families of functions from the considered classes. We observe that studying of equi-families can be reduced to the exploration of a single orbit map with values in a product space. We consider the closure of equi-families with respect to the topology of pointwise convergence. Finally, we investigate functions $f\colon X\times Y\to Z$, for metric spaces $X,Y,Z$, with sections that are equi-continuous, equi-Baire 1 or have equi-generalized Lebesgue property with respect to measurable sets of class $\alpha$. In particular, we generalize a result of Grande.
Key words and phrases:
Borel 1 functions, Lebesgue property, weakly separated functions, equi-Baire 1 families, topology of pointwise convergence, functions in two variables of Borel class $\alpha$
2020 Mathematics Subject Classification:
54H05, 26A21
1. Introduction
Let $f$ be a function between two metric spaces $X$ and $Y$. We say that $f$ is Baire 1 if $f$ is the limit of a pointwise convergent sequence of continuous functions $f_{n}\colon X\to Y$, $n\in\mathbb{N}$. We say that $f$ is Borel 1 if it is $F_{\sigma}$-measurable, that is, its preimage of every open set is $F_{\sigma}$. Note that every Baire 1 function is Borel 1, but the converse does not hold in general; see [21, 24B]. However, if $Y=\mathbb{R}$ or $Y$ is a Banach space, then the Baire 1 functions and the Borel 1 functions coincide; see [21] and also [13] and [28, Remark 5.6] for more general results. Note that several authors (see, e.g. [21], [26], [27]) use the name Baire 1 for $F_{\sigma}$-measurable functions which can sometimes lead to confusion.
Many nice characterizations of Baire 1 (or Borel 1) functions are known. Let us mention some of them. The classical theorem of Baire states that, if $X$ is complete and $Y$ is separable, then $f$ is Borel 1 if and only if $f\upharpoonright P$ has a point of continuity for every non-empty closed set $P\subseteq X$ (see [24]); for some generalizations, see [23]). This last condition is called the Point of Continuity Property (PCP).
Lee, Tang, and Zhao [27] obtained an interesting $\varepsilon$-$\delta$ characterization of Borel 1 functions between two Polish spaces. The respective condition characterizing Borel 1 functions was called LTZ-property in [4]. The result of [27] was later generalized in [11] to the case when $X$ and $Y$ are separable spaces. The authors of [11] used an old Lebesgue theorem [25] in their proof. Although the Lebesgue result was concerned with the case $X=Y=\mathbb{R}$, it remains true if $X$ is arbitrary and $Y$ is separable [4, Thm 2.1]. The respective condition stated in Lebesgue’s result was referred to in [4] as the Lebesgue property.
Note that Baire 1 real-valued functions and their regular subclasses have important applications in classical theory of real functions (see [7]) and functional analysis [16]. Independently, some other interesting classes of maps related to Baire 1 functions have been studied, also in a general setting where $X$ is a topological space. Namely, Koumoullis [23] considered fragmentable maps and functions with PCP. Bouziad [6] investigated weakly separated maps, which coincide with functions with LTZ-property if $X$ is a metric space. He also used cliquish and fragmentable maps, as well as those having PCP.
In Section 2 we collect the above-mentioned classes of maps. We introduce a new notion — the generalized Lebesgue property. We establish inclusions between these families. First, this is done when $X$ is a topological space. Some of the inclusions can be reversed under additional assumptions in $X$, while others are appropriate, as shown by examples. Second, we consider the case where both spaces $X$ and $Y$ are metric spaces.
Our work is, in some sense, a continuation of [5] and [4], which were inspired by the intriguing notion of equi-Baire 1 families of functions, proposed independently by Lecomte [26] and Alikhani-Koopeai [2]. This notion plays a role in dynamical systems similar to that of equi-continuous families (see [3]). We observed that other types of equi-families of functions can be of interest (see [14], [15] and [4], where equi-PCP property and equi-Lebesgue property were considered). We collect such families in Section 3 and present a simple tool which shows that the study of equi-families can be reduced to the analysis of a single orbit map with values in the product space. Finally, we solve two problems from [5] concerning functions of two variables with sections being equi-continuous or equi-Borel measurable of class $\alpha$, and we obtain a generalization of the result of Grande [15].
2. Properties related to being a Borel 1 map
We consider several collections of functions from a topological space $X$ into a metric space $Y$ that are related to Borel 1 functions and which turned out to be important when studying families of equi-Baire 1 functions. We establish inclusions between them and show examples when inclusions are proper. If $X$ is a metric space, some of these families are equal under or without additional assumptions on $X$.
2.1. Relations between various families of functions
Definition 2.1.
Let $X$ be a topological space, $(Y,d)$ be a metric space and $f\colon X\to Y$ be a function. We say that $f$
  • is weakly separated, if
    • (WS)
      for every $\varepsilon>0$ there exists a neighborhood assignment $\{V_{x}\}_{x\in X}$ such that for all $x,y\in X$
      $(x,y)\in V_{y}\times V_{x}\,\,\,\Longrightarrow\,\,\,d(f(x),f(y))<\varepsilon;$
      (A family $\{V_{x}\}_{x\in X}$ of open sets in $X$ is called a neighborhood assignment whenever $x\in V_{x}$ for each $x\in X$.)
  • is fragmentable, if
    • (Frag)
      for every $\varepsilon>0$ and every closed set $\emptyset\neq F\subseteq X$ there exists an open set $U$ such that $U\cap F\neq\emptyset$ and ${\rm diam}f(U\cap F)<\varepsilon$;
  • is cliquish, if
    • (Cliq)
      for every $\varepsilon>0$ and every open set $\emptyset\neq U\subseteq X$ there exists an open set $\emptyset\neq O\subseteq U$ such that ${\rm diam}f(O)<\varepsilon$;
    (Clearly, $f$ is fragmentable if and only if $f\restriction_{F}$ is cliquish for each nonempty closed set $F$.)
  • has the Lebesgue property, if
    • (LP)
      for every $\varepsilon>0$ there is a cover $(X_{n})_{n\in\mathbb{N}}$ of $X$ consisting of closed sets such that ${\rm diam}f(X_{n})\leq\varepsilon$ for all $n\in\mathbb{N}$;
  • has the generalized Lebesgue property, if
    • (GLP)
      for every $\varepsilon>0$ there is a $\sigma$-discrete family $\mathcal{A}_{\varepsilon}$ consisting of closed subsets of $X$ such that $X=\bigcup\mathcal{A}_{\varepsilon}$ and ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}_{\varepsilon}$.
    (Recall that a family $\mathcal{A}$ of subsets of topological space $X$ is discrete, if every point $x\in X$ has a neighborhood $U$ which intersects at most one set from the family $\mathcal{A}$. It is worth noting that the union of an arbitrary discrete family of closed sets is also a closed set. A family $\mathcal{A}$ is called $\sigma$-discrete, if it is a countable union of discrete families.)
  • has the point of continuity property (PCP), if
    • (PCP)
      for every closed set $\emptyset\neq F\subseteq X$ the restriction $f\restriction_{F}$ has a point of continuity;
  • is pointwise discontinuous (PWD), if
    • (PWD)
      for every open set $\emptyset\neq U\subseteq X$ the restriction $f\restriction_{U}$ has a point of continuity;
    (In other words, $f$ is pointwise discontinuous whenever the set of continuity points of $f$ is dense in its domain.)
  • is Borel 1, if
    • (Borel 1)
      the preimage $f^{-1}(V)$ of any open set $V\subseteq Y$ is an $F_{\sigma}$-set in $X$.
Theorem 2.2.
For any topological space $X$ and a metric space $Y$ the following relations hold:
WSCliqPWDFragGLPBorel 1LPPCPhBBYY sepBhB
Here (“hB”) “B” near an arrow means that $X$ is (hereditarily) Baire, “$Y$ sep” means that $Y$ is separable.
Proof.
Implications (PCP) $\Rightarrow$ (Frag) and (Frag) $\Rightarrow$ (PCP) for hereditarily Baire spaces were proved by Koumoullis in [23, Theorem 2.3]. Further, Bouziad [6] showed that (Frag) $\Rightarrow$ (WS). Implication (Frag) $\Rightarrow$ (Cliq) follows easily from the definitions. Actually, Lebesgue proved that (LP) $\Rightarrow$ (Borel 1) (see [24, p. 375] and [4]). Lebesgue proved also that (Borel 1) $\Rightarrow$ (LP) (see also [24, p. 375] and [4, Theorem 2.1]) for a separable space $Y$. Implication (Cliq) $\Rightarrow$ (PWD) was proved in [9]. Since every countable family $\mathcal{A}$ is $\sigma$-discrete, we have (LP) $\Rightarrow$ (GLP).
Let us show that (GLP) $\Rightarrow$ (WS). Assume $f\colon X\to Y$ has the generalized Lebesgue property and fix $\varepsilon>0$. There exists a closed cover $\mathcal{A}$ of $X$ such that ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}$, and $\mathcal{A}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$ where $\mathcal{A}_{n}$ is discrete for each $n\in\mathbb{N}$. Note that for all $n\in\mathbb{N}$ and $x\in\bigcup\mathcal{A}_{n}$ there exists a unique $F^{n}_{x}\in\mathcal{A}_{n}$ such that $x\in F^{n}_{x}$. Let $X_{n}:=\bigcup\mathcal{A}_{n}$ for each $n\in\mathbb{N}$. For every $x\in X$ we put
$n(x):=\min\{n:x\in X_{n}\}\textup{\ and\ }V_{x}:=X\setminus\bigcup_{n<n(x)}X_{n}\setminus\bigcup\{F\in\mathcal{A}_{n(x)}\colon x\notin F\}.$
Then $\{V_{x}\}_{x\in X}$ is a neighborhood assignment such that $V_{x}\cap X_{n(x)}\subseteq F^{n(x)}_{x}$ for each $x\in X$. Assume $(x,y)\in V_{y}\times V_{x}$. Then $x\in X_{n(x)}$ and $x\not\in X_{n}$ for all $n<n(y)$, consequently $n(x)\geq n(y)$. Similarly, $n(y)\geq n(x)$. Hence, $n(x)=n(y)$. Since $x\in V_{y}\subseteq F^{n(y)}_{y}=F^{n(x)}_{y}$, we have $F^{n(x)}_{y}=F^{n(x)}_{x}$. Thus $d(f(x),f(y))\leq{\rm diam}f(F^{n(y)}_{y})\leq\varepsilon$.
We show that (GLP) $\Rightarrow$ (Borel 1). Let $f\colon X\to Y$ have the generalized Lebesgue property. For every $n\in\mathbb{N}$ we choose a family $\mathcal{A}$ of subsets of $X$ such that $\mathcal{A}_{n}=\bigcup_{k\in\mathbb{N}}\mathcal{A}_{n,k}$, where each $\mathcal{A}_{n,k}$ is a discrete family consisting of closed sets, $\mathcal{A}_{n}$ is a covering of $X$ and ${\rm diam}f(A)<1/n$ for each $A\in\mathcal{A}_{n}$. Denote $\mathcal{B}:=\bigcup_{n,k\in\mathbb{N}}\mathcal{A}_{n,k}$. Notice that for every subfamily $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ the union $\bigcup\mathcal{B}^{\prime}$ of all members in $\mathcal{B}^{\prime}$ is an $F_{\sigma}$-set, since every $\mathcal{A}_{k,n}$ is a discrete family of closed sets. It is easy to see that
$f^{-1}(V)=\bigcup\{B\in\mathcal{B}:B\subseteq f^{-1}(V)\}\textup{\ for any open set\ }V\subseteq Y.$
Hence, $f^{-1}(V)$ is an $F_{\sigma}$-set in $X$. Therefore, $f$ is Borel 1.
Finally, we show that (GLP) $\Rightarrow$ (Cliq) if $X$ is a Baire space (see also [17, proof of Theorem 3.8]). Let $f\colon X\to Y$ have the generalized Lebesgue property, $\varepsilon>0$ and $\emptyset\neq U\subseteq X$ be an open set. Similarly as in the proof of “(GLP) $\Rightarrow$ (WS)” there exists a closed cover $\mathcal{A}_{\varepsilon}$ of $X$ such that ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}$, and $\mathcal{A}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$, where $\mathcal{A}_{n}$ is discrete for each $n\in\mathbb{N}$. Note that for all $n\in\mathbb{N}$ and $x\in\bigcup\mathcal{A}_{n}$ there exists a unique $F^{n}_{x}\in\mathcal{A}_{n}$ such that $x\in F^{n}_{x}$. Let $X_{n}:=\bigcup\mathcal{A}_{n}$ for each $n\in\mathbb{N}$. Note that $(X_{n})_{n\in\mathbb{N}}$ is a cover of $X$ consisting of closed sets. Since $U=\bigcup_{n\in\mathbb{N}}(X_{n}\cap U)$ and $U$ is a Baire space, there is a nonempty open set $V\subseteq U$ and $n_{0}\in\mathbb{N}$ such that $V\subseteq X_{n_{0}}\cap U$. Fix $x\in V$. Since $\mathcal{A}_{n_{0}}$ is discrete, there exists a nonempty open set $V^{\prime}$ such that $V^{\prime}\cap X_{n_{0}}=V^{\prime}\cap F^{x}_{n_{0}}$. Then ${\rm diam}f(V\cap V^{\prime})\leq\varepsilon$.
[…]
Definition 2.3.
We say that a function $f\colon X\to Y$ between a metric space $(X,\varrho)$ and a metric space $(Y,d)$ has the Lee-Tang-Zhao (LTZ) property, if
  • (LTZ)
    for every $\varepsilon>0$ there is a positive function $\delta_{\varepsilon}\colon X\to\mathbb{R}_{+}$ such that for all $x,y\in X$
    $\varrho(x,y)<\min\left\{\delta_{\varepsilon}(x),\delta_{\varepsilon}(y)\right\}\,\,\,\Longrightarrow\,\,\,d(f(x),f(y))<\varepsilon.$
Theorem 2.4.
For any metric spaces $X$ and $Y$ the following relations hold
WSCliqPWDFragGLPBorel 1LPPCPhBBYY sepBhBhBXX sephBLTZ
Here (“hB”) “B” near an arrow means that $X$ is (hereditarily) Baire, “$X$ sep” (“$Y$ sep”) means that $X$ (respectively, $Y$) is separable, dashed line means that $X$ is metric and solid line means that $X$ is arbitrary topological (solid part of the diagram is covered by Theorem 2.2).
Proof.
Assume (where it is indicated in the diagram) that $X$ is hereditarily Baire. Condition (WS) $\Leftrightarrow$ (LTZ) is an obvious observation; it was explicitly stated in introduction to [6].
Implications (Frag) $\Rightarrow$ (Borel 1) and (Borel 1) $\Rightarrow$ (PCP) were proved by Koumoullis [23, Theorem 4.12].
From the proof of Lee, Tang and Zhao [27] it follows that (LTZ) $\Rightarrow$ (PCP) for a hereditarily Baire $X$.
Implication (Borel 1) $\Rightarrow$ (WS) for a hereditarily Baire space $X$ was proved in [19, Theorem 12].
Implications (GLP) $\Rightarrow$ (LTZ) and (LTZ) $\Rightarrow$ (GLP) were proved in [19] (see Lemma 4 and Lemma 5, respectively).
To see (GLP) $\Rightarrow$ (LP) for separable metric spaces $X$ observe that, if $X$ is metric and separable, then a $\sigma$-discrete family must be countable. (Recall that a straightforward proof of (LTZ) $\Rightarrow$ (LP) for separable $X$ was given in [11, Theorem 5]). ∎
Remark 2.5.
Implication (WS) $\Rightarrow$ (Borel 1) for a hereditarily Baire semistratifiable space $X$ was also proved by Bouziad [6, Theorem 4.5] (in fact, he proved that (WS) $\Rightarrow$ (GLP)).
The equivalence (Frag) $\Leftrightarrow$ (GLP) was proved in [20, Theorem B] for a hereditarily Baire perfect paracompact space $X$.
Implication (GLP) $\Rightarrow$ (LP) works also for a collectionwise normal separable space. Every paracompact space is collectionwise normal. (For the definition of a collectionwise normal space, see [10, p. 305].)
Corollary 2.6.
Let $X$ be a hereditarily Baire metric space and $Y$ is a metric space. Then
(GLP) $\Leftrightarrow$ (WS) $\Leftrightarrow$ (LTZ) $\Leftrightarrow$ (PCP) $\Leftrightarrow$ (Borel 1) $\Leftrightarrow$ (Frag).
Moreover, if either $X$ or $Y$ is separable, then
(LP) $\Leftrightarrow$ (GLP) $\Leftrightarrow$ (WS) $\Leftrightarrow$ (LTZ) $\Leftrightarrow$ (PCP) $\Leftrightarrow$ (Borel 1) $\Leftrightarrow$ (Frag).
2.2. Examples
In this section we give counterexamples for implications not included in Theorem 2.2 and 2.4. Case (Cliq) $\not\Rightarrow$ (PWD) and (Frag) $\not\Rightarrow$ (PCP) are covered by Proposition 2.12 and Proposition 2.13, respectively.
Example 2.7.
(WS) $\not\Rightarrow$ (Frag) even if $X$ is metric.
Proof.
Let $X=\mathbb{Q}$ and $A,B\subseteq X$ be disjoint sets such that $\overline{A}=\overline{B}=X$. Then the function $f\colon X\to\mathbb{R}$, $f=\chi_{A}$, is Borel 1 and weakly separated, but ${\rm diam}f(U)=1$ for every open nonempty set $U\subseteq X$. Hence, $f$ is not fragmentable.
Since every function on $\mathbb{Q}$ is Borel 1, we get (Borel 1) $\not\Rightarrow$ (Frag). ∎
Example 2.8.
(Cliq) $\not\Rightarrow$ (Borel 1) even if $X$ is metric and hereditarily Baire.
Proof.
Let $X=Y=\mathbb{R}$ and $C\subseteq X$ be the Cantor set. Denote by $A$ the set of all endpoints of $C$ and put $f=\chi_{A}$. It is easy to check that $f\colon X\to Y$ is cliquish, but $f\restriction_{C}$ is not Borel 1 because $A$ is not a $G_{\delta}$-set in $C$. Hence, $f\colon X\to Y$ is not Borel 1.
As a corollary we obtain (Cliq) $\not\Rightarrow$ (Frag). ∎
Example 2.9.
(WS) $\not\Rightarrow$ (Borel 1) even if $X$ is hereditarily Baire and $Y$ is separable.
Proof.
Let $X=\mathbb{S}$ be the Sorgenfrey line and $Y=\mathbb{R}$. Consider $f\colon X\to Y$, $f=\chi_{\mathbb{Q}}$. Then $f$ is weakly separated, since it is easy to see that every function defined on $\mathbb{S}$ is weakly separated. But $f$ is not Borel 1 because $\mathbb{Q}$ is not a $G_{\delta}$-set in $X$.
Consequently, (WS) $\not\Rightarrow$ (GLP). Moreover, (WS) $\not\Rightarrow$ (PCP) if $X$ is not metrizable. ∎
Example 2.10 (Martin’s Axiom and the negation of the Continuum Hypothesis).
(Borel 1) $\not\Rightarrow$ (LP) even if $X$ is metric and separable.
Proof.
Under Martin’s Axiom and the negation of the Continuum Hypothesis, there exists an uncountable set $X\subseteq\mathbb{R}$ such that each subset of $X$ is $F_{\sigma}$ in $X$; see [8, Theorem 9.6]. Assume that $Y=X$ is endowed with the discrete metric, i.e. $d(x,y)=1$ if $x\neq y$ and $d(x,y)=0$ if $x=y$. Then the function $f\colon X\to Y$, $f(x)=x$, is Borel 1. Assume that $f$ has the Lebesgue property. For $\varepsilon=\tfrac{1}{2}$ we choose a closed covering $\{F_{n}:n\in\mathbb{N}\}$ of $X$ such that ${\rm diam}f(F_{n})<\tfrac{1}{2}$. Then ${\rm diam}f(F_{n})=0$ which implies that $f(F_{n})=\{y_{n}\}$ for every $n\in\mathbb{N}$. Since the sequence $(F_{n})$ covers $X$, it follows that $Y$ is countable, a contradiction. ∎
Note that, if $X$ is a nonseparable metric space, then the identity function $\operatorname{Id}\colon X\to X$ is Borel 1 and it does not have (LP); see [4, Proposition 2.1].
2.3. More examples: Baire spaces and cliquish/fragmentable functions
In this section, we show in the form of a characterization that some implications in Theorem 2.2 are not reversible.
Lemma 2.11.
Let $X$ be a topological space. If $X$ is not a Baire space, then there are a nonempty open set $U\subseteq X$ and a fragmentable function $f\colon X\to\mathbb{R}$ such that $f$ has no point of continuity in $U$ and $f(X)\subseteq\{1/n:n\in\mathbb{N}\}$. Moreover, if $X$ is perfectly normal, then $f$ has the Lebesgue property.
Proof.
If $X$ is not a Baire space, then there is a nonempty open set $U$ in $X$ such that $U$ is of the first Baire category in $X$. Let $\{F_{n}:n\in\mathbb{N}\}$ be a sequence of relatively closed sets in $U$, which are nowhere dense in $X$ and such that $F_{n}\subseteq F_{n+1}$ for every $n\in\mathbb{N}$, and $U=\bigcup_{n\in\mathbb{N}}F_{n}$.
Define a function $f\colon X\to\mathbb{R}$ by the formula:
$f(x)=\left\{\begin{array}[]{ll}1&\textup{if\ }x\in F_{1}\cup(X\setminus U),\\ \frac{1}{n}&\textup{if\ }x\in F_{n}\setminus F_{n-1},n\in\mathbb{N},n>1.\end{array}\right$.
It is easy to verify that there is no point of continuity of the function $f$ in the open set $U$. (If $n>1$, then $f^{-1}(\{1/n\})=f^{-1}((1/(n+1),1/(n-1))=F_{n}\setminus F_{n-1}$ is a nowhere dense set in $X$, so it contains no nonempty open subset. Also $f^{-1}(\{1\})\cap U=f^{-1}((1/2,3/2))\cap U=F_{1}$ is a nowhere dense set in $X$.
We prove that $f$ is fragmentable. Let $\varepsilon>0$ and $F$ be a nonempty closed set in $X$. If $F\subseteq X\setminus U$, then ${\rm diam}f(X\cap F)=0<\varepsilon$. Suppose $F\cap U\neq\emptyset$. If $F\cap U\cap F_{n}\neq\emptyset$ for infinitely many $n\in\mathbb{N}$, let $k\in\mathbb{N}$ be such that $2/k<\varepsilon$. Put $U\setminus F_{k-1}. GG$ is a nonempty open set in $X$ and ${\rm diam}f(G\cap F)<2/k<\epsilon$. Suppose there is $n\in\mathbb{N}$ such that $F\cap U\subseteq F_{n}$. Without loss of generality we can suppose that $n\in\mathbb{N}$ is minimal such that $F\cap U\subseteq F_{n}$. If $n=1$, then ${\rm diam}f(F\cap U)=0<\varepsilon$. If $n>1$, the set $U\setminus F_{n-1}$ is a nonempty open set in $X$ and ${\rm diam}f(U\setminus F_{n-1})=0<\varepsilon$.
To see that $f$ has the Lebesgue property it is enough to observe that for a perfectly normal $X$ the set $F_{n}\setminus F_{n-1}$ is $F_{\sigma}$ for each $n$. ∎
Proposition 2.12.
Let $X$ be a topological space. The following are equivalent:
  1. (1)
    $X$ is a Baire space;
  2. (2)
    every cliquish $f\colon X\to\mathbb{R}$ is pointwise discontinuous.
Moreover, if $X$ is perfectly normal, the above conditions are equivalent to:
  1. (3)
    every $f\colon X\to\mathbb{R}$ with the Lebesgue property is pointwise discontinuous.
Proof.
[…]
Proposition 2.12 has its “hereditary” variant:
Proposition 2.13.
Let $X$ be a topological space. The following are equivalent:
  1. (1)
    $X$ is a hereditarily Baire space;
  2. (2)
    if $f\colon X\to\mathbb{R}$ is fragmentable, then $f$ has (PCP).
Moreover, if $X$ is perfectly normal, the above conditions are equivalent to:
  1. (3)
    if $f\colon X\to\mathbb{R}$ has the Lebesgue property, then $f$ has (PCP).
Proof.
[…]
We show that $g$ is fragmentable. Let $\varepsilon>0$ and $C$ be a closed nonempty subset of $X$. If $C\cap(X\setminus F)\neq\emptyset$, then ${\rm diam}g(C\cap(X\setminus F))=0<\varepsilon$.
Suppose now that $C\subseteq F$. Since the function $f\colon F\to\mathbb{R}$ is fragmentable there is a nonempty open set $U$ in $F$ such that $C\cap U\neq\emptyset$ and ${\rm diam}f(C\cap U)<\varepsilon$. There is an open set $G$ in $X$ such that $U=G\cap F$. Thus we have $G\cap C=G\cap C\cap F$ and ${\rm diam}g(G\cap C)={\rm diam}f(C\cap U)<\varepsilon$.
[…]
Theorem 2.4 and Proposition 2.13 imply the following result.
Corollary 2.14.
Let $X$ be a metric space. The following are equivalent:
  1. (1)
    $X$ is a hereditarily Baire space;
  2. (2)
    if $f\colon X\to\mathbb{R}$ is Borel 1, then $f$ has (PCP).
Remark 2.15.
Observe that, in the additional assertions of Lemma 2.11 and Proposition 2.13, we can only assume that every open set in $X$ is $F_{\sigma}$ instead of supposing that $X$ is perfectly normal.
3. Equi-families of functions
For collections of functions investigated in Section 2, we study the respective subfamilies that fulfill the defining condition in common, the so-called equi-families with the adequate property. If we consider the respective orbit map with values in the product space, its property is equivalent to the respective equi-property for a given subfamily.
We then present solutions of two problems from [5]. Finally, we generalize the result of Grande [15] showing a sufficient condition for $f\colon X\times Y\to Z$ to be a Borel function of class $\alpha$, in the language of $x$-sections and $y$-sections of $f$.
3.1. Equi-families and the orbit function
The notion of an equi-Baire 1 family introduced by [26] was a starting point in the previous papers [5] and [4]. Let $(X,\rho)$ and $(Y,d)$ be metric spaces and $\mathcal{F}\subseteq Y^{X}$. Then $\mathcal{F}$ is called equi-Baire 1 if
  • for every $\varepsilon>0$ there is a positive function $\delta_{\varepsilon}\colon X\to\mathbb{R}_{+}$ such that for all $x,y\in X$ and all $f\in\mathcal{F}$
    $\varrho(x,y)<\min\left\{\delta_{\varepsilon}(x),\delta_{\varepsilon}(y)\right\}\,\,\,\Longrightarrow\,\,\,d(f(x),f(y))<\varepsilon.$
In other words, a modification of (LTZ) property holds where $\delta_{\varepsilon}$ is chosen the same for every $f\in\mathcal{F}$.
In the same spirit, we can modify several properties considered in Section 2 for a single function $f$ from a topological space $X$ to a metric space $(Y,d)$, and we obtain the respective equi-properties for a family $\mathcal{F}\subseteq Y^{X}$. We mention that equi-Lebesgue and equi-cliquish families were considered in [4].
Definition 3.1.
Let $\mathcal{F}\subseteq Y^{X}$ be a family of functions between a topological space $X$ and a metric space $(Y,d)$. We say that
  • $\mathcal{F}$ is equi-weakly separated, if for every $\varepsilon>0$ there exists a neighborhood assignment $\{V_{x}\}_{x\in X}$ such that for all $x,y\in X$ and all $f\in\mathcal{F}$
    $(x,y)\in V_{y}\times V_{x}\,\,\,\Longrightarrow\,\,\,d(f(x),f(y))<\varepsilon;$
  • $\mathcal{F}$ is equi-fragmentable, if for every $\varepsilon>0$ and every closed set $\emptyset\neq F\subseteq X$ there exists an open set $U$ such that $U\cap F\neq\emptyset$ and ${\rm diam}f(U\cap F)<\varepsilon$ and for all $f\in\mathcal{F}$;
  • $\mathcal{F}$ is equi-cliquish, if for every $\varepsilon>0$ and every open set $\emptyset\neq U\subseteq X$ there exists an open set $\emptyset\neq O\subseteq U$ such that ${\rm diam}f(O)<\varepsilon$ for all $f\in\mathcal{F}$;
  • $\mathcal{F}$ is equi-Lebesgue, if for every $\varepsilon>0$ there is a cover $(X_{n})_{n\in\mathbb{N}}$ of $X$ consisting of closed sets such that ${\rm diam}f(X_{n})\leq\varepsilon$ for all $n\in\mathbb{N}$ and all $f\in\mathcal{F}$;
  • $\mathcal{F}$ is equi-GLP, if for every $\varepsilon>0$ there is a $\sigma$-discrete family $\mathcal{A}_{\varepsilon}$ consisting of closed subsets of $X$ such that $X=\bigcup\mathcal{A}_{\varepsilon}$ and ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}_{\varepsilon}$ and all $f\in\mathcal{F}$;
  • $\mathcal{F}$ has the point of equicontinuity property (PECP), if for every closed set $\emptyset\neq F\subseteq X$ the family $\{f\restriction_{F}\colon f\in\mathcal{F}\}$ has a point of equicontinuity.
Remark 3.2.
Clearly, if $X$ is a metric space in the above definition, the notions of an equi-weakly separated family and an equi-Baire 1 family coincide.
Let $X$ be a topological space, and $(Y,d)$ be a bounded metric space. For a family $\mathscr{F}\subseteq Y^{X}$ of functions, we denote by
$f^{\sharp}_{\mathscr{F}}(x)=(f(x))_{f\in\mathscr{F}}$
the orbit function $f^{\sharp}_{\mathscr{F}}\colon X\to Y^{T}$, where $T=|\mathscr{F}|$ (see e.g. [1, Sec. 4]). Assume that $Z=Y^{T}$ is equipped with the supremum metric
$\varrho(z_{1},z_{2})=\sup\limits_{t\in T}d(z_{1}(t),z_{2}(t)).$
It is easy to see that the following observation is valid. This is an extended version of Proposition 1 from [18].
Proposition 3.3.
Let $X$ be a topological space, and $(Y,d)$ be a bounded metric space. Then
  1. (1)
    $\mathscr{F}$ is equi-continuous at $x\in X$ (on $X$) if and only if $f^{\sharp}_{\mathscr{F}}\colon X\to(Z,\varrho)$ is continuous at $x$ (on $X$);
  2. (2)
    $\mathscr{F}$ is equi-weakly separated if and only if $f^{\sharp}_{\mathscr{F}}\colon X\to(Z,\varrho)$ is weakly separated;
  3. (3)
    if $X$ is metric, then $\mathscr{F}$ is equi-Baire 1 if and only if $f^{\sharp}_{\mathscr{F}}\colon X\to(Z,\varrho)$ has LTZ-property;
  4. (4)
    $\mathscr{F}$ is equi-fragmentable if and only if $f^{\sharp}_{\mathscr{F}}\colon X\to(Z,\varrho)$ is fragmentable;
  5. (5)
    $\mathscr{F}$ is equi-cliquish if and only if $f^{\sharp}_{\mathscr{F}}\colon X\to(Z,\varrho)$ is cliquish;
  6. (6)
    $\mathscr{F}$ is equi-Lebesgue if and only if $f^{\sharp}_{\mathscr{F}}\colon X\to(Z,\varrho)$ has the Lebesgue property;
  7. (7)
    $\mathscr{F}$ is equi-GLP if and only if $f^{\sharp}_{\mathscr{F}}\colon X\to(Z,\varrho)$ has the generalized Lebesgue property;
  8. (8)
    $\mathscr{F}$ has (PECP) if and only if $f^{\sharp}_{\mathscr{F}}\colon X\to(Z,\varrho)$ has (PCP).
We will show some applications of this proposition.
Remark 3.4.
Let $X$ be a topological space and $(Y,d)$ be a metric space. Put $d^{*}:=\min\{d,1\}$. Let $\mathcal{F}\subseteq Y^{X}$. Clearly, $\mathcal{F}$ has any equi-property from Definition 3.1 using the metric $d$ if and only if it has the respective equi-property using the metric $d^{*}$. Similarly, if $X$ is a metric space, $\mathcal{F}$ is equi-Baire 1 using the metric $d$ if and only if it has equi-Baire 1 property using the metric $d^{*}$.
Proposition 3.5.
Let $X$ and $Y$ be metric spaces and $\mathcal{F}\subseteq Y^{X}$.
  1. (1)
    If $\mathcal{F}$ has (PECP), then $\mathcal{F}$ is equi-Baire 1.
  2. (2)
    Assume that $X$ is hereditary Baire. Then $\mathcal{F}$ is equi-Baire 1 if and only if $\mathcal{F}$ has (PECP).
Proof.
It is enough to apply Theorem 2.4, Proposition 3.3 and Remark 3.4. Note that (2) generalizes the result of [26, Proposition 32]. ∎
The following proposition extends Proposition 6.4 from [4].
Proposition 3.6.
Let $X$ be a Baire topological space, $Y$ be a metric space, and $\mathcal{F}\subseteq Y^{X}$.
  1. (1)
    If $\mathcal{F}$ is equi-GLP, then $\mathcal{F}$ is equi-cliquish.
  2. (2)
    If $X$ is a metric space and $\mathcal{F}$ is equi-Baire 1, then $\mathcal{F}$ is equi-cliquish.
Proof.
Assertion (1) follows from Theorem 2.2, Proposition 3.3 and Remark 3.4. Assertion (2) follows from Theorem 2.4, Proposition 3.3 and Remark 3.4. ∎
For a family $\mathcal{F}$ of functions from a topological space $X$ to a metric space $Y$ we denote by $EC(\mathcal{F})$ the set of equi-continuity points of $\mathcal{F}$. It is known that this set is $G_{\delta}$ since $EC=\bigcap_{n\in\mathbb{N}}O(1/n)$ where
$O(\varepsilon)=\bigcup\{U\colon U\subseteq X,\;U\text{ nonempty open, diam}f(U)<\varepsilon\text{ for all }f\in\mathcal{F}\}$
is open for each $\varepsilon>0$; see [12].
We can generalize Proposition 6.3 from [4] as follows.
Proposition 3.7.
Let $X$ be a Baire topological space, $Y$ be a metric space, and $\mathcal{F}\subseteq Y^{X}$.
  1. (1)
    If $\mathcal{F}$ is equi-cliquish, then $EC(\mathcal{F})$ is a dense $G_{\delta}$ set. Thus, if $\mathcal{F}$ is equi-GLP, then $EC(\mathcal{F})$ is a dense $G_{\delta}$ set.
  2. (2)
    If $X$ is a metric space and $\mathcal{F}$ is equi-Baire 1, then $EC(\mathcal{F})$ is a dense $G_{\delta}$ set.
Proof.
Assertion (1) follows from Theorem 2.2, Proposition 3.3 and Remark 3.4. Assertion (2) follows from Theorem 2.4, Proposition 3.3 and Remark 3.4. ∎
3.2. Closures in the topology of pointwise convergence
The following result is classical.
Proposition 3.8.
([22], Theorem 14, Chapter 7) Let $X$ be a topological space, $(Y,d)$ be a metric space and $\mathcal{F}$ be a family of functions from $X$ to $Y$. If $\mathcal{F}$ is equicontinuous at $x$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is also equicontinuous at $x$.
We can state similar results for other equi-families.
Proposition 3.9.
([26], Proposition 35) Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces. If $\mathcal{F}$ is an equi-Baire 1 family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-Baire 1 too.
Proposition 3.10.
([4], Proposition 4.1) Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces. If $\mathcal{F}$ is an equi-Lebesgue family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-Lebesgue too.
Proposition 3.11.
Let $X$ be a topological space, and let $(Y,d)$ be a metric space. If $\mathcal{F}$ is an equi-weakly separated family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is also equi-weakly separated.
Proof.
Let $\varepsilon>0$. Since $\mathcal{F}$ is an equi-weakly separated family of functions from $X$ to $Y$, there exists a neighbourhood assignment $(V_{x})_{x\in X}$ such that for every $x,y\in X$ and $f\in\mathcal{F}$
$(x,y)\in V_{y}\times V_{x}\textup{\ implies\ }d(f(x),f(y))\leq\varepsilon/3.$
Then for $\varepsilon$ and the neighbourhood assignment $(V_{x})_{x\in X}$ we have: for every $x,y\in X$ and for every $g\in\overline{\mathcal{F}}$
$(x,y)\in V_{y}\times V_{x}\textup{\ implies\ }d(f(x),f(y))\leq\varepsilon.$
Let $(x,y)\in V_{y}\times V_{x}$. If $g\in\mathcal{F}$, then $d(g(x),g(y))\leq\varepsilon/3$. Let $g\in\overline{\mathcal{F}}\setminus\mathcal{F}$. There is a function $f\in\mathcal{F}$ such that $d(f(x),g(x))<\varepsilon/3$ and $d(f(y),g(y)))<\varepsilon/3$. Then we have
$d(g(x),g(y))\leq d(g(x),f(x))+d(f(x),f(y))+d(f(y),g(y))\leq\varepsilon.$
Using similar ideas, we can also prove the following implications.
Proposition 3.12.
Let $X$ be a topological space and $(Y,d)$ be a metric space.
  1. (1)
    If $\mathcal{F}$ is an equi-Lebesgue family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-Lebesgue too.
  2. (2)
    If $\mathcal{F}$ is an equi-GLP family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is also equi-GLP.
  3. (3)
    If $\mathcal{F}$ is an equi-cliquish family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-cliquish too.
  4. (4)
    If $\mathcal{F}$ is an equi-fragmentable family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-fragmentable too.
  5. (5)
    If $\mathcal{F}$ has (PECP), then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence has (PECP) too.
3.3. Applications to functions defined on $X\times Y$
We are going to solve two problems from the earlier paper [5].
Problem 3.13.
[5, Prob. 3.7] Suppose that $\mathcal{F}\subseteq\mathbb{R}^{X\times Y}$ is a family of separately equi-continuous functions defined on the product of Polish spaces $X$ and $Y$, i.e.
  • $\{f(x,\cdot)\colon f\in\mathcal{F}\}$ is equi-continuous for each $x\in X$, and
  • $\{f(\cdot,y)\colon f\in\mathcal{F}\}$ is equi-continuous for each $y\in Y$.
Is $\mathcal{F}$ equi-Baire 1?
Let us give a positive solution in a general setting. Bouziad [6, Proposition 4.3] proved the following fact. If $X,Y$ are topological spaces and $Z$ is a metric space, then every separately continuous function $f\colon X\times Y\to Z$ is weakly separated. Now, we obtain the following corollary which solves the problem.
Corollary 3.14.
Let $X,Y$ be topological spaces and $Z$ be a metric space. Let $\mathcal{F}\subseteq Z^{X\times Y}$ be a family of separately equi-continuous functions defined on $X\times Y$. Then $\mathcal{F}$ is equi-weakly separated.
Proof.
Consider the orbit function $f^{\sharp}_{\mathscr{F}}\colon X\times Y\to Z^{T}$ (where $T:=|\mathscr{F}|$) as in Proposition 3.3. Note that it is separately continuous. By the Bouziad result, $f^{\sharp}_{\mathscr{F}}$ is weakly separated and we infer that $\mathscr{F}$} is equi-weakly separated by Proposition 3.3. ∎
Let us state the next problem [5, Prob. 3.11] in a strictly formal version as follows.
Problem 3.15.
Suppose that $X$ and $Y$ are Polish and for each $n\in\mathbb{N}$ we have $f_{n}\colon X\times Y\to\mathbb{R}$. Moreover, suppose that:
  • the family $\{f_{n}(x,\cdot)\colon n\in\mathbb{N}\}$ is equi-Baire 1 for each $x\in X$, and
  • the family $\{f_{n}(\cdot,y)\colon n\in\mathbb{N}\}$ is equi-Baire 1 for each $y\in Y$.
Does there exist a comeager $G_{\delta}$ set $A\subseteq X$ such that $\{f_{n}|_{A\times Y}\colon n\in\mathbb{N}\}$ is jointly equi-Baire 1?
Remark 3.16.
Using Proposition 3.3(6) for the families:
  • $\mathcal{F}:=\{f_{n}\colon n\in\mathbb{N}\}$,
  • $\mathcal{F}^{1}_{x}:=\{f_{n}(x,\cdot)\colon n\in\mathbb{N}\}$ for each $x\in X$,
  • $\mathcal{F}^{2}_{y}:=\{f_{n}(\cdot,y)\colon n\in\mathbb{N}\}$ for each $y\in Y$,
we can see that the above problem is equivalent to the question whether for each separately Baire 1 function $f^{\#}\colon X\times Y\to\mathbb{R}^{\mathbb{N}}$ there exists a comeager $G_{\delta}$ set $A\subseteq X$ such that $f^{\#}|_{A\times Y}$ is jointly Baire 1.
Let us show a negative solution of Problem 3.15. Grande noticed that there exists a function $f\colon\mathbb{R}^{2}\to\mathbb{R}$ that has all $x$-sections and all $y$-sections of class Baire 1 but $f$ is not Lebesgue measurable (or it does not have the Baire property); see e.g. [15, Remark 2]. One can use this fact to obtain the respective counterexample for Problem 3.15. However, we propose a different simple construction. Let $B\subseteq\mathbb{R}$ be a Bernstein nonmeasurable set. It is known that for any uncountable Borel set $A\subseteq\mathbb{R}$ the intersection $B\cap A$ is not a Borel set. Let $g\colon\mathbb{R}^{2}\to\mathbb{R}$ be the characteristic function of the set $B^{*}=\{(x,x)\colon x\in B\}$. Then $g$ is separately Baire 1. Suppose $A\subseteq\mathbb{R}$ is a nonmeager $G_{\delta}$ set such that $g\restriction_{A\times\mathbb{R}}$ is Baire 1. Then $A\cap B$ should be Borel, a contradiction. Similarly, the function $f\colon\mathbb{R}^{2}\to\mathbb{R}^{\mathbb{N}}$ given by $f=(g,g,\dots)$ is separately Baire 1 and $f\restriction_{A\times\mathbb{R}}$ is not jointly Baire 1 for every nonmeager $G_{\delta}$ set $A\subseteq\mathbb{R}$. This by Remark 3.16 leads to the negative solution of Problem 3.15.
Grande proved in [15, Thm 6] the following interesting fact on real functions of two variables. If $f\colon\mathbb{R}^{2}\to\mathbb{R}$ is a function with all $x$-sections that form a family having (PECP) and all $y$-sections of Borel class $\alpha$, then $f$ is of Borel class $\alpha$. We have observed that this result with the same proof remains valid if $f\colon X\times Y\to Z$ where $X,Y,Z$ are metric spaces and $Y$ is separable. Below, we propose a further extension of this theorem.
Definition 3.17.
We say that a function $f\colon X\to Y$ between a topological space $X$ and a metric space $Y$ has $\alpha$-generalized Lebesgue property, for $1\leq\alpha<\omega_{1}$, if
  • ($\alpha$-GLP)
    for every $\varepsilon>0$ there is a $\sigma$-discrete family $\mathcal{A}_{\varepsilon}$ in $X$ consisting of Borel sets of additive class $\alpha$ such that $X=\bigcup\mathcal{A}_{\varepsilon}$ and ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}_{\varepsilon}$.
Remark 3.18.
It is easy to see that for an isolated ordinal $\alpha=\beta+1$ we may assume that all elements from $\mathcal{A}_{\varepsilon}$ belongs to the multiplicative class $\beta$. Hence, Definition 3.17 for $\alpha=1$ is consistent with the definition of the generalized Lebesgue property.
If $\alpha$ is limit, then we may assume that all elements of $\mathcal{A}$ belongs to multiplicative classes less than $\alpha$. Since every set of multiplicative class less than $\alpha$ in a metric space is ambiguous of class $\alpha$, one can substitute the word “additive” by “ambiguous” in Definition 3.17.
Definition 3.19.
Let $\mathscr{F}\subseteq Y^{X}$ be a family of functions between $X$ and $Y$, and $1\leq\alpha<\omega_{1}$. We say that $\mathscr{F}$ is equi-$\alpha$-GLP, if ($\alpha$-GLP) holds for every $f\in\mathcal{F}$, i.e. for every $\varepsilon>0$ there is a $\sigma$-discrete family $\mathcal{A}_{\varepsilon}$ in $X$ consisting of Borel sets of additive class $\alpha$ such that $X=\bigcup\mathcal{A}_{\varepsilon}$ and ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}_{\varepsilon}$ and $f\in\mathcal{F}$.
If $f\colon X\times Y\to Z$, then for $x\in X$ and $y\in Y$ by $f^{x}$ and $f_{y}$ we denote $x$-section $f^{x}\colon Y\to Z$ and $y$-section $f_{y}\colon X\to Z$ of the function $f$, respectively.
Let $\mathcal{A}$ and $\mathcal{B}$ be families of sets. We write $\mathcal{A}\prec\mathcal{B}$ if for every set $A\in\mathcal{A}$ there exists a set $B\in\mathcal{B}$ such that $A\subseteq B$.
Theorem 3.20.
Let $X,Y$ and $(Z,d)$ be metric spaces, $1\leq\alpha<\omega_{1}$ and let a function $f\colon X\times Y\to Z$ satisfy the following properties:
  1. (1)
    the family $\mathscr{F}_{X}=\{f^{x}:x\in X\}$ is equi-$\alpha$-GLP,
  2. (2)
    for every $y\in Y$ the section $f_{y}\colon X\to Z$ belongs to the Borel class $\alpha$.
Then $f\colon X\times Y\to Z$ belongs to the Borel class $\alpha$.
Proof.
We will show that for every $\varepsilon>0$ there exists a function $g\colon X\times Y\to Z$ of Borel class $\alpha$ such that
$d(f(p),g(p))\leq\varepsilon$ for all $p\in X\times Y$.
Fix $\varepsilon>0$. Since $\mathscr{F}_{X}$ is equi-$\alpha$-GLP, there exists a $\sigma$-discrete family $\mathcal{A}$ in $Y$ consisting of sets of additive and multiplicative class $\alpha$ simultaneously such that
$Y=\bigcup\mathcal{A}$  and  ${\rm diam}\,f^{x}(A)\leq\varepsilon$
for all $A\in\mathcal{A}$ and for all $x\in X$. Let $\mathcal{A}=\bigcup\limits_{n}\mathcal{A}_{n}$, where each family $\mathcal{A}_{n}$ is discrete in $Y$. We put
$\mathcal{B}_{n}=(A\setminus\bigcup\limits_{k<n}\bigcup\mathcal{A}_{k}:A\in\mathcal{A}_{n}),\,\,n\in\mathbb{N},$
$B_{n}=\bigcup\{B:B\in\mathcal{B}_{n}\}$ and  $\mathcal{B}=\bigcup\limits_{n=1}^{\infty}\mathcal{B}_{n}.$
Then $\mathcal{B}$ is a $\sigma$-discrete family of sets of ambiguous class $\alpha$ with $\mathcal{B}\prec\mathcal{A}$ and $\bigcup\mathcal{B}=Y$. Moreover, $B^{\prime}\cap B^{\prime\prime}=\emptyset$ for all $B^{\prime}\in\mathcal{B}_{k}$, $B^{\prime\prime}\in\mathcal{B}_{n}$ and distinct $k,n\in\mathbb{N}$. Notice that every set $B_{n}$ belongs to the additive and multiplicative class $\alpha$ simultaneously as the union of a discrete family of sets of the same classes in metric space $Y$.
Take an arbitrary $y_{B,n}\in B$ for every $B\in\mathcal{B}_{n}$, $n\in\mathbb{N}$. For each $y\in Y$ we put $n(y)=\min\{n\in\mathbb{N}:y\in B_{n}\}$. Now let $(x,y)\in X\times Y$. Since the family $\mathcal{B}_{n(y)}$ is disjoint, there exists a unique set $B(y)\in\mathcal{B}_{n(y)}$ such that $y\in B(y)$. We define
$g(x,y)=f(x,y_{B(y),n(y)}).$
Then for every point $(x,y)\in X\times Y$ we have
$d(f(x,y),g(x,y))=d(f(x,y),f(x,y_{B(y),n(y)}))\leq{\rm diam}\,f^{x}(B(y)),$
since $y,y_{B(y),n(y)}\in B(y)$. It follows from the property $\mathcal{B}\prec\mathcal{A}$ that ${\rm diam}\,f^{x}(B(y))\leq{\rm diam}\,f^{x}(A)$ for some $A\in\mathcal{A}$. Hence,
$d(f(x,y),g(x,y))\leq\varepsilon.$
It remains to prove that $g\colon X\times Y\to Z$ belongs to the Borel class $\alpha$. Fix $n\in\mathbb{N}$ and let $g_{n}=g\restriction_{X\times B_{n}}$. It follows from the definition of $g$ and assumption (2) of the theorem, that $g_{n,B}=g_{n}\restriction_{X\times B}$ belongs to the Borel class $\alpha$ for every $B\in\mathcal{B}_{n}$. Then for every open set $W\subseteq Z$ we have that the preimage
$g_{n}^{-1}(W)=\bigcup\{g_{n,B}^{-1}(W):B\in\mathcal{B}_{n}\}$
is of the additive class $\alpha$ in $X\times B_{n}$ as a union of a discrete family of sets of the same class. Now, since $\{X\times B_{n}:n\in\mathbb{N}\}$ is a countable covering of $X\times Y$ by sets of ambiguous class $\alpha$ and $g\restriction_{X\times B_{n}}$ is Borel $\alpha$, we conclude that $g\colon X\times Y\to Z$ belongs to the Borel class $\alpha$.
Finally, let us remember that the family of functions of Borel class $\alpha$ between metric spaces is closed under uniform limits. ∎
By Proposition 3.3 and Theorem 2.4:
$\textrm{(PECP)}\Rightarrow\textrm{(equi-fragmentable)}\Rightarrow\textrm{(equi-Baire 1)}\Rightarrow\textrm{(equi-GLP)}\Rightarrow\textrm{(equi-$\alpha$-GLP)},$
so the following corollary is a generalization of Grande’s result.
Corollary 3.21.
Let $X$, $Y$ and $Z$ be metric spaces and $f\colon X\times Y\to Z$ satisfies the following properties
  1. (1)
    the family $\mathscr{F}_{X}=\{f^{x}:x\in X\}$ has (PECP),
  2. (2)
    for every $y\in Y$ the section $f_{y}\colon X\to Z$ belongs to the Borel class $\alpha$.
Then $f\colon X\times Y\to Z$ belongs to the Borel class $\alpha$.
Ackonwledgement. We would like to thank Tomek Natkaniec for showing us Theorem 6 in Grande’s article [15] which is relevant to our consideration of Problem 3.15.